# Conformalized Interval Arithmetic with Symmetric Calibration

### v1

@@ Appendix C Analysis of CIA on Overlapping Subsets
Figure 4. The figure illustrates the positive correlation between the subsets overlap ratio $\delta$ and the coverage gap (the difference between the coverage probability and the desired $1-\alpha$ coverage).
@@
Overall, the experimental results support the theoretical findings and demonstrate the effectiveness of CIA in providing valid confidence intervals for overlapping subsets.
Appendix D Summary of Methods and Results
D.1. Methods
D.1.1. CIA Variants
  1. (1)
    (8) $s^{\textrm{cal}}_{k}:=\left|\sum_{i\in S^{\textrm{cal}}_{k}}y_{i}-\widehat{y}_{i}\right|,k\in[K].$
  2. (2)
    (9) $s^{\textrm{cal}}_{k}=\max\left\{\sum_{i\in S^{\textrm{cal}}_{k}}\widehat{y}_{i}^{\alpha/2}-y_{i},\sum_{i\in S^{\textrm{cal}}_{k}}y_{i}-\widehat{y}^{1-\alpha/2}_{i}\right\},k\in[K].$
  3. (3)
  4. (4)
D.1.2. Baseline Using Conformal Prediction
  1. (1)
  2. (2)
  3. (3)
  4. (4)
D.1.3. Baseline Using Normal Approximation
  1. (1)
  2. (2)
D.2. Results
We present the comparison results of the aforementioned methods across various datasets as follows. The results demonstrate the performance of each method in terms of coverage probability and set size.
Figure 5. Complete results for prediction sets. All CIA variants achieve $1-\alpha$ coverage, with CIA (CQR) Stratified showing optimal efficiency.
Table 1. Coverage and Size Results for Different Methods Across Datasets at $\alpha=0.1$
@@ D.2. Results
Table 2. Coverage and Size Results for Different Methods Across Datasets at $\alpha=0.05$
@@
Table 3. Coverage and Size Results for Different Methods Across Datasets at $\alpha=0.01$

### v2

@@ Appendix D Analysis of CIA on Overlapping Subsets
Figure 4: The figure illustrates the positive correlation between the subsets overlap ratio $\delta$ and the coverage gap (the difference between the coverage probability and the desired $1-\alpha$ coverage).
@@
Overall, the experimental results support the theoretical findings and demonstrate the effectiveness of CIA in providing valid confidence intervals for overlapping subsets, with a reasonable trade-off between coverage probability and interval width.
Figure 5: Complete results for prediction sets. All CIA variants achieve $1-\alpha$ coverage, with CIA (CQR) Stratified showing optimal efficiency.
Table 1: Coverage and Size Results for Different Methods Across Datasets at $\alpha=0.1$
@@
Table 2: Coverage and Size Results for Different Methods Across Datasets at $\alpha=0.05$
@@
Table 3: Coverage and Size Results for Different Methods Across Datasets at $\alpha=0.01$